2. Compute the direct labor cost variance, including its rate and efficiency variances. Standard Cost Actual Cost $ $ 0 0 0 03. Compute the overhead controllable and volume variances. Controllable Variance Actual overhead Budgeted overhead ControllabCheck my work 1 ! Required information 12.5 [The following information applies to the questions displayed below.] points TricDuring the current quarter, the company operated at 90% of capacity and produced 50,400 units of product; actual direct laborRequired: 1. Compute the direct materials cost variance, including its price and quantity variances. Actual Cost Standard Cos
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1499000 $6.30 $9,443,700.00 1499000 $5.10 $7,644,900.00 1512000 $5.10 $7,711,200.00
$1,798,800.00 U $66,300.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $1,798,800.00 U
Direct material quantity variance $66,300.00 F
Direct material cost variance $1,732,500.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
299400 $12.50 $3,742,500.00 299400 $15.00 $4,491,000.00 302400 $15.00 $4,536,000.00
$748,500.00 F $45,000.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $748,500.00 F
Direct Labor Efficiency variance $45,000.00 F
Direct labor cost variance $793,500.00 F

Solution 3:

Controllable Variance
Actual overhead $5,043,200.00
Budgeted overhead $5,073,600.00
Controllable variance $30,400.00 F
Fixed overhead volume variance
Budgeted fixed overhead $2,956,800.00
Fixed overhead cost applied $3,326,400.00
Fixed overhead volume variance $369,600.00 F